Determine the pH of each of the following solutions.

(a) 0.416 M hypobromous acid (weak acid with Ka = 2.5e-09).

pH = ___________

(b) 0.644 M ascorbic acid (weak acid with Ka = 8e-05).

pH = ___________

(c) 0.703 M pyridine (weak base with Kb = 1.7e-09).

pH = ___________
Answer #1

0.415 Ka= acid Hypobromous 27.5 xlo-9 or weak acid [H3O+] = kałHOBr] 7.5x10-9x0.416 = 3.2249 x10-5 PH= -log (3.22ug X10-5) PH